OpenStudy (anonymous):

I give a medal to the person that helps!!: What is the length of a side of a cube with a volume of 343 cm3?

ganeshie8 (ganeshie8):

remember the volume of cube formula ?

OpenStudy (anonymous):

yea

ganeshie8 (ganeshie8):

good, equate that formula to 343

ganeshie8 (ganeshie8):

cuz you're given volume is 343

ganeshie8 (ganeshie8):

wat do u get ?

OpenStudy (anonymous):

a=7 cm

OpenStudy (anonymous):

Oh! Alright thanks guys I figured it out :)

ganeshie8 (ganeshie8):

good :) \(\large Volume = s^3 = 343\) so, \(\large s = \sqrt[3]{343} = 7\)